1.Potassium nitrate decomposes to potassium nitrite and oxygen. 2KNO3(s)→2KNO2(s)+O2(g)

What volume, in liters, of O2 can be produced from the decomposition of 54.0 g of KNO3 at 33 ∘C and 1.11 atm ?

Express your answer with the appropriate units.

2,On a climb up Mt. Whitney, the atmospheric pressure drops to 507 mmHg .

A.What is the pressure in atmospheres?Express the pressure in atmospheres to three significant figures.

B.What is the pressure in torrs? Express the pressure in torrs to three significant figures.

C.What is the pressure in inHg?Express the pressure in inches of mercury to three significant figures.

D.What is the pressure in Pa?Express the pressure in pascals to three significant figures.
Answer #1

1)

Molar mass of KNO3,

MM = 1*MM(K) + 1*MM(N) + 3*MM(O)

= 1*39.1 + 1*14.01 + 3*16.0

= 101.11 g/mol

mass of KNO3 = 54 g

mol of KNO3 = (mass)/(molar mass)

= 54/1.011*10^2

= 0.5341 mol

According to balanced equation

mol of O2 formed = (1/2)* moles of KNO3

= (1/2)*0.5341

= 0.267 mol

Given:

P = 1.11 atm

n = 0.267 mol

T = 33.0 oC

= (33.0+273) K

= 306 K

use:

P * V = n*R*T

1.11 atm * V = 0.267 mol* 0.08206 atm.L/mol.K * 306 K

V = 6.04 L

Answer: 6.04 L

Only 1 question at a time please